HUBBLE SPACE TELESCOPE DAILY REPORT #1095 PERIOD COVERED: 0700 EST Daily Status Report as of 074/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxies 3C330.0 and 3C228.0 as part of a continuum survey and to study the galaxies' morphology. Four images were received and displayed. All four were reported to look okay. 1.2 Completed WF/PC-2 5565 (WF/PC-2 Cycle 4: Photometric Calibration Monitor: Fields) The WF/PC-2 was used to make photometric calibration images by taking images of the large globular cluster Omega Centauri (aka, NGC-5139) to help determine astrometric and optical stability of the WF/PC-2. The OSS monitored the observations, and assessed the data quality. A total of ten observations were monitored. No anomalies were noted. 1.3 Completed WF/PC-2 5612 (SMOV WF/PC-2 Decontamination) The WF/PC-2 was used to do decontamination operations. Observations of standard A+81D266, and K-Spot observations were also taken. A total of eight observations were monitored and displayed. The internal and external observations were all nominal. 1.4 Completed WF/PC-2 4709 (WF/PC-2 Internal Monitor) The WF/PC camera was used to make a series of routine ultraviolet flat calibration images to monitor the instrument's performance. The observations were monitored in the OSS, and the data quality was assessed. 26 observations were received and displayed. One observations was reported to have been missing one group from a partial STR dump. The version of the file from the dump with the rest of the data appeared to be intact, but the first group could not be displayed. 1.5 Completed FGS 4884 (High Seed Astrometry: A Search for Astrometric Companions to Low Mass Stars) The FGSs were used to make observations of Bernard's star to look for astrometric companions (i.e., a Jupiter-sized companion). The first four observations were monitored, and appeared nominal. The remainder of the observations occurred during LOS, and will be analyzed later. 1.6 Completed Three Sets of WF/PC-2 5090 (Large Area: Multicolor Survey: Cycle 4) The WF/PC-2 was used to make deep exposure images of fields #1416 and #1418 as part of a large area, high galactic latitude, multicolor sky survey to learn more about the distribution of galaxies, their morphology, distribution in magnitude and redshift, and evolution of galactic luminosity function and colors. Eight images were received and displayed. The first six were reported to be nominal. The last two were reported to have had recentering events occur during the observations. 1.7 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and for contamination. A total of ten images were taken. Six were received and displayed, and four remain on the science tape recorder. No anomalies were noted. 1.8 Completed WF/PC-2 5562 (WF/PC-2 Cycle 4 Calibration: Internal Monitor 3: Darks) The WF/PC-2 was used to obtain several 30 minute long dark exposures to provide ongoing calibration of the CCD dark current and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled :17 Successful :17 2.3 Significant HSTARs: None 2.4 Operations Notes: SMOV activities performed last week included: flat fields calibration, decontamination monitoring, and the completion of fine alignment and focus operations with the final SMOV adjustment of the articulated fold mirrors for WF/PC-2; and the repeat of FOS/FGS fine alignment of the FOS/Red large apertures. SMOV functions remaining to be conducted are: FOS/FGS fine alignment of small apertures, FOS plate scale, GHRS early release observation (ERO), and a repeat of the FOS ERO to correct a target acquisition/pointing problem. With the exception of the last item, all SMOV activities are scheduled for completion before or during the last week of March. For the last 24 hours, there have been no reported problems with any SMOV operation or with any HST subsystem. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Three Sets of WF/PC-2 5090 (Large Area/Multicolor Survey: Cycle 4) 3.2 Complete GHRS/2 5549 (GHRS/2 Side 2 Dark Noise Monitor: Cycle 4) 3.3 Complete GHRS/2 5546 (GHRS Side 2 Threshold Adjustment Test: Cycle 4) 3.4 Complete Two Sets of 5570 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Large Set) 3.5 Complete WF/PC-2 5476 (Continuum Snapshots of 3RC Radio Galaxies) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-